OpenStudy (maddyiscoolio):

The formula below gives A, the area of a circle with radius r. A(r)=3.14r to the second power To the nearest hundredth what is the value of A(6)? A.)18.84 B.)37.68 C.)354.95 D.)113.04

OpenStudy (aaronandyson):

Replace the "r" with 6?

OpenStudy (maddyiscoolio):

but then its like 0.523333 and that is not what its asking..

OpenStudy (mathstudent55):

\(\large A(r) = 3.14r^2\) is the formula of the area of a circle, where r is the radius. \(\large A(6)\) means the area of circle whose radius is 6. To find the area, substitute r with 6 in the formula. \(\large A(6) = 3.14 \times 6^2\)

OpenStudy (mathstudent55):

First, find \(6^2\). Remember that an exponent (whole number 2 or more) means to multiply the base a number of times. With exponent 2, you multiply the base by itself. \(6^2 = 6\times 6\) Once you do that multiplication, then multiply by 3.14

OpenStudy (mathstudent55):

What do you get for the area?

OpenStudy (mathmale):

Once you obtain a result for the area, it's up to YOU to round off this result to the nearest hundredth.